# Methods and Trends in Detecting Generated Images: A Comprehensive Review

### v1

@@ 2.1.8. Classification and Explainable Identification: Bird and Lotfi (2024)
2.1.9. Self-Contrastive Learning on ImagiNet: Boychev et al. (2024)
Boychev et al. (Boychev and Cholakov, 2024) introduce a large-scale dataset, ImagiNet, consisting of 200,000 real and synthetic images, as detailed in the datasets section. The detection framework consists of two stages: pretraining with Self-Contrastive Learning (SelfCon) and a calibration step. During the pretraining phase, the ResNet-50 (He et al., 2016) backbone, initialized with ImageNet weights, is paired with a sub-network that projects intermediate feature maps into a shared latent space. This setup produces two output embeddings per input image, facilitating contrastive learning. The SelfCon loss is defined as:
@@
Table 1. Evaluation of Cross-Family Generators, Cross-Category, and Cross-Scene Generalization in Detection Models from Spatial Domain Analysis Category
@@
Nataraj et al. (Nataraj et al., 2019) X X X
Wang et al. (Wang et al., 2020b) X X
Goebel et al. (Goebel et al., 2020) X X
Li et al. (Li et al., 2022) X X X
Lorenz et al. (Lorenz et al., 2023) X
Baraheem et al. (Baraheem and Nguyen, 2023) X X
Wang et al. (Wang et al., 2023a)
Bird & Lotfi (Bird and Lotfi, 2024) X X
Boychev et al. (Boychev and Cholakov, 2024) X
@@
Figure 3. Taxonomy of Vision-Language Models (VLMs) categorized by learning paradigms.
[…]
Diagram illustrating the types of VLMs
2.2. Multimodal Vision-Language Methods
@@
Multimodal vision-language methods leverage vision-language models (VLMs) (see Fig. 3 for the types of VLMs) trained on large-scale image-text datasets to detect generated images by aligning visual and textual features. These approaches detect inconsistencies in synthetic images through cross-modal embeddings, enabling effective detection across a wide range of generative methods. Models such as CLIP (Radford et al., 2021) which belongs to the contrastive learning (Chen et al., 2020) category (see Fig. 3 for categorization) are commonly employed in transfer learning setups, with novel adaptations to distinguish real images from generated ones. These methods demonstrate robust generalization and adaptability to various generative techniques and contexts, as will be discussed in the following paragraphs.
2.2.1. Universal Fake Image Detector by Ojha (2023)
@@
Ojha et al. (Ojha et al., 2023) identified that existing fake image detectors struggle with generalization, often misclassifying images from unseen generative models as real. This limitation arises from classifiers being asymmetrically tuned to detect artifacts specific to training data. To address this, the authors propose constructing a meaningful feature space using CLIP:ViT (Dosovitskiy, 2020; Radford et al., 2021), trained on 400 million image-text pairs. They employ two approaches: (a) mapping training images to CLIP’s final layer to create feature representations, then during inference, classifying images based on cosine distance to the nearest neighbor in real and fake feature spaces, and (b) augmenting CLIP with a linear layer for binary classification. Both methods demonstrated generalization, effectively detecting synthetic images from state-of-the-art generative models.
2.2.2. Language-Guided Synthetic Image Detection by Wu (2023)
@@
Wu et al. (Wu et al., 2023) propose a language-guided approach for detecting synthetic images by integrating image-text contrastive learning in a VLM. They reformulate the detection task as an identification problem, determining whether a query image aligns with an anchor set of text-labeled images. This method enhances generalization by aligning visual features with carefully designed textual labels such as ”Real Photo,” ”Real Painting,” ”Synthetic Photo,” and ”Synthetic Painting.” The authors found these labels more effective than simpler ”real” or ”fake” categories, as they account for differences in image types, such as camera-captured versus digitally created content. The proposed LASTED framework encodes images using a ResNet-50x64 vision encoder and text labels using a transformer-based text encoder. During training, both encoders generate visual and textual representations, $e_{v}$ and $e_{t}$, respectively. A contrastive loss aligns these features, ensuring that matched pairs have higher similarity scores than unmatched ones:
@@
2.2.3. GenDet: Good Generalizations by Zhu (2023)
Zhu et al. (Zhu et al., 2023) address the challenge of detecting synthetic images from unseen generators, which existing methods struggle to classify due to limited generalization. The authors propose GenDet, a detection model composed of two key components: Teacher-Student Discrepancy-Aware Learning and Generalized Feature Augmentation. These components are trained through an adversarial framework to improve generalization to both seen and unseen generators. The model employs a feature extractor $E$, based on CLIP:ViT (Radford et al., 2021), to extract features from input images. The Teacher-Student Discrepancy-Aware Learning is designed to: a) Reduce the difference in output between a teacher network ($N_{t}$) and a student network ($N_{s}$) for real images, b) Amplify this difference for fake images to enhance detection. The discrepancy losses are defined as:
@@ 2.2.3. GenDet: Good Generalizations by Zhu (2023)
2.2.4. Mixture of Low-Rank Experts: Liu 2024
Liu et al. (Liu et al., 2024b) propose a transferable AI-generated image detection model utilizing CLIP-ViT as the backbone with parameter-efficient fine-tuning. The method modifies the MLP layers of the last three ViT-B/32 blocks through a Mixture of Low-Rank Experts (MoLE), integrating both shared and separate Low-Rank Adapters (LoRAs). Shared LoRAs capture common feature representations across datasets, while separate LoRAs specialize in diverse generative patterns. A trainable gating mechanism dynamically assigns input tokens to appropriate experts, with a load-balancing loss ensuring uniform expert utilization. The model freezes most CLIP parameters, adapting only LoRA modules and a new MLP classification head with a sigmoid activation. The forward operation in each MLP block is expressed as:
@@ 2.2.4. Mixture of Low-Rank Experts: Liu 2024
2.2.5. MERGING A MIXTURE OF HYPER LORAS: HYPERDET by Cao 2024
Cao et al. (Cao et al., 2024) propose a generalizable synthetic image detection framework, HyperDet, leveraging the large pretrained multimodal model, CLIP: ViT-L/14, as the backbone, similar to Liu et al. (Liu et al., 2024b), but with few notable innovations. Unlike Liu et al., the authors introduce a novel grouping of Spatial Rich Model (SRM) filters into five distinct groups to generate multiple filtered views of input images, capturing varying levels of pixel artifacts and features. Along this, HyperDet employs Hyper LoRAs, a hypernetwork-based approach that generates Low-Rank Adaptation (LoRA) weights for fine-tuning the CLIP model. These LoRA weights are computed using three types of embeddings: task embeddings, layer embeddings, and position embeddings. The outputs of these LoRA experts are merged to form a unified representation for classification, effectively integrating shared and specific knowledge for generalizable feature extraction. During training, HyperDet fine-tunes the last eight fully connected layers of the CLIP: ViT-L/14 model, along with the newly introduced Hyper LoRAs modules. To address imbalanced optimization, the framework employs a composite binary cross-entropy loss function, incorporating both original and filtered views of the images. This design obtains robust performance in detecting synthetic images across diverse generative models and datasets.
2.2.6. Forgery-Aware Adaptive Transformer (FatFormer) by Liu: 2024
Liu et al. (Liu et al., 2024a) introduce FatFormer, a generalizable synthetic image detection framework utilizing the pretrained CLIP model inspired by the work of Ojha et al. (Ojha et al., 2023). The authors address the limitations of freezing CLIP’s layers, which hinders the generalization of forgery detection. FatFormer integrates two modules: the Forgery-Aware Adapter (FAA) and Language-Guided Alignment (LGA), for effective adaptation of CLIP’s features. The FAA module extracts forgery artifacts from both image and frequency domains. The Image Forgery Extractor applies lightweight convolution layers to capture low-level artifacts, while the Frequency Forgery Extractor employs Discrete Wavelet Transform (DWT) and grouped attention mechanisms to dynamically aggregate multi-band frequency clues. The final adapted feature representation at each ViT stage is defined as:
@@ 2.2.6. Forgery-Aware Adaptive Transformer (FatFormer) by Liu: 2024
2.2.7. Raising the Bar with CLIP: By Cozzolino 2024
Cozzolino et al. (Cozzolino et al., 2024) implement the CLIP: ViT-L/14 pretrained VLM, similar to the approaches in (Cao et al., 2024; Liu et al., 2024b; Koutlis and Papadopoulos, 2025), for detecting synthetic images with a straightforward yet impactful adjustment. The authors propose generating synthetic images by feeding real-image captions to text-to-image models and then extracting feature vectors for both real and synthetic images using CLIP’s image encoder. Specifically, feature vectors are obtained from the second-to-last layer of the ViT module:
@@ 2.2.7. Raising the Bar with CLIP: By Cozzolino 2024
2.2.8. Representation from Encoder-Decoder for Image Detection by Koutlis: 2025
Koutlis and Papadopoulos (Koutlis and Papadopoulos, 2025) propose Representations from Intermediate Encoder-Blocks (RINE) to improve synthetic image detection by extracting low-level features from multiple layers of CLIP’s Vision Transformer (ViT). The method captures both low- and high-level visual semantics by concatenating CLS tokens from each intermediate transformer block into a comprehensive feature representation. The extracted CLS tokens from each block $Z_{l}$ are aggregated:

### v2

@@ 2.1.7 Classification and Explainable Identification: Bird and Lotfi (2024)
2.1.8 Self-Contrastive Learning on ImagiNet: Boychev et al. (2024)
Boychev et al. Boychev and Cholakov (2024) introduce a large-scale dataset, ImagiNet, consisting of 200,000 real and synthetic images, as detailed in the datasets section. The detection framework consists of two stages: pretraining with Self-Contrastive Learning (SelfCon) and a calibration step. During the pretraining phase, the ResNet-50 He et al. (2016) backbone, initialized with ImageNet weights, is paired with a sub-network that projects intermediate feature maps into a shared latent space. This setup produces two output embeddings per input image, facilitating contrastive learning. The SelfCon loss is defined as:
@@
Comparative Analysis of Spatial Domain Analysis Methods
Spatial-domain forensic methods share the unified goal of identifying AI-generated imagery through pixel-level and localized artifact analysis, yet they differ in their architectural principles and scope of generalization. Early CNN-based detectors, such as Wang et al. Wang et al. (2020b), revealed that discriminative spatial artifacts are shared across different CNN generators, demonstrating cross-model generalization using a standard ResNet-50 backbone trained solely on ProGAN data. However, these early detectors were limited in interpretability and reliance on dataset-specific features. Goebel et al. Goebel et al. (2020) addressed these issues by introducing co-occurrence matrix–based representations that capture pixel dependencies across multiple orientations, enabling both detection and attribution within a unified framework. To overcome the poor cross-domain transferability of such handcrafted representations, Li et al. Li et al. (2022) proposed GASE-Net, emphasizing representation learning and artifact similarity estimation between domain prototypes and consult images, which improved robustness to post-processing and domain shifts. Lorenz et al. Lorenz et al. (2023) further expanded generalization beyond GANs by introducing AdaptedMultiLID, which models intrinsic feature-space dimensionality rather than explicit pixel statistics, addressing the prior limitation of model specificity. Baraheem et al. Baraheem and Nguyen (2023) contributed the RSI benchmark and demonstrated that transfer learning on large-scale diverse GAN datasets with EfficientNet and CAM visualizations enhances explainability and model generalization. To address the gap in detecting diffusion-generated content, Wang et al. Wang et al. (2023a) introduced DIRE, which integrates spatial-domain analysis with diffusion-based reconstruction, achieving higher performance on unseen diffusion architectures. More recently, Bird and Lotfi Bird and Lotfi (2024) and Boychev et al. Boychev and Cholakov (2024) tackled the remaining limitations of explainability and training efficiency by incorporating Grad-CAM interpretability and self-contrastive learning (SelfCon), respectively, offering improved transparency and zero-shot generalization. Collectively, these studies illustrate a clear methodological evolution—from handcrafted co-occurrence features and CNN-based spatial detection, to representation and reconstruction learning, and finally to contrastive and explainable frameworks. The trajectory underscores the field’s movement from model-specific and dataset-dependent detection toward unified, interpretable, and generalizable spatial-domain analysis for AI-generated image forensics.
Table 1: Evaluation of Cross-Family Generators, Cross-Category, and Cross-Scene Generalization in Detection Models from Spatial Domain Analysis Category
@@
Wang et al. Wang et al. (2020b) X X
Goebel et al. Goebel et al. (2020) X X
Li et al. Li et al. (2022) X X X
Lorenz et al. Lorenz et al. (2023) X
Baraheem et al. Baraheem and Nguyen (2023) X X
Wang et al. Wang et al. (2023a)
Bird & Lotfi Bird and Lotfi (2024) X X
Boychev et al. Boychev and Cholakov (2024) X
@@
Figure 3: Taxonomy of Vision-Language Models (VLMs) categorized by learning paradigms.
2.2 Multi-modal Vision-Language Methods (and/or Multi-modal Large Language Models)
@@
This section encompasses two closely related but distinct sub-categories of forensic detection methods: Multimodal Vision-Language Models (VLMs) and Multimodal Large Language Models (MLLMs). The first category leverages vision-language models (VLMs) (see Fig. 3 for representative architectures) trained on large-scale image–text datasets to detect generated images by aligning visual and textual features. These approaches primarily depend on a visual encoder and typically lack explicit reasoning capabilities. They detect inconsistencies in synthetic images through cross-modal embeddings, enabling robust detection across a broad range of generative models. Representative examples include CLIP Radford et al. (2021), which belongs to the contrastive learning paradigm Chen et al. (2020) (see Fig. 3 for categorization), and its numerous transfer-learning adaptations that distinguish real images from generated ones. The second category, developed more recently, predominantly employs Multimodal Large Language Models (MLLMs) that jointly utilize visual and textual inputs to provide not only detection but also interpretive reasoning explaining the model’s decision. These methods commonly adopt recent VLM or MLLM backbones such as Qwen Bai et al. (2025), LLaMA, LLaVA Liu et al. (2023), and GPT Achiam et al. (2023), achieving enhanced explainability and alignment with human-understandable reasoning.
2.2.1 Universal Fake Image Detector by Ojha (2023)
@@
Ojha et al. Ojha et al. (2023) identified that existing fake image detectors struggle with generalization, often misclassifying images from unseen generative models as real. This limitation arises from classifiers being asymmetrically tuned to detect artifacts specific to training data. To address this, the authors propose constructing a meaningful feature space using CLIP:ViT Dosovitskiy (2020); Radford et al. (2021), trained on 400 million image-text pairs. They employ two approaches: (a) mapping training images to CLIP’s final layer to create feature representations, then during inference, classifying images based on cosine distance to the nearest neighbor in real and fake feature spaces, and (b) augmenting CLIP with a linear layer for binary classification. Both methods demonstrated generalization, effectively detecting synthetic images from state-of-the-art generative models.
2.2.2 Language-Guided Synthetic Image Detection by Wu (2023)
@@
Wu et al. Wu et al. (2023) propose a language-guided approach for detecting synthetic images by integrating image-text contrastive learning in a VLM. They reformulate the detection task as an identification problem, determining whether a query image aligns with an anchor set of text-labeled images. This method enhances generalization by aligning visual features with carefully designed textual labels such as "Real Photo," "Real Painting," "Synthetic Photo," and "Synthetic Painting." The authors found these labels more effective than simpler "real" or "fake" categories, as they account for differences in image types, such as camera-captured versus digitally created content. The proposed LASTED framework encodes images using a ResNet-50x64 vision encoder and text labels using a transformer-based text encoder. During training, both encoders generate visual and textual representations, $e_{v}$ and $e_{t}$, respectively. A contrastive loss aligns these features, ensuring that matched pairs have higher similarity scores than unmatched ones:
@@
2.2.3 GenDet: Good Generalizations by Zhu (2023)
Zhu et al. Zhu et al. (2023) address the challenge of detecting synthetic images from unseen generators, which existing methods struggle to classify due to limited generalization. The authors propose GenDet, a detection model composed of two key components: Teacher-Student Discrepancy-Aware Learning and Generalized Feature Augmentation. These components are trained through an adversarial framework to improve generalization to both seen and unseen generators. The model employs a feature extractor $E$, based on CLIP:ViT Radford et al. (2021), to extract features from input images. The Teacher-Student Discrepancy-Aware Learning is designed to: a) Reduce the difference in output between a teacher network ($N_{t}$) and a student network ($N_{s}$) for real images, b) Amplify this difference for fake images to enhance detection. The discrepancy losses are defined as:
@@ 2.2.3 GenDet: Good Generalizations by Zhu (2023)
2.2.4 Mixture of Low-Rank Experts: Liu 2024
Liu et al. Liu et al. (2024b) propose a transferable AI-generated image detection model utilizing CLIP-ViT as the backbone with parameter-efficient fine-tuning. The method modifies the MLP layers of the last three ViT-B/32 blocks through a Mixture of Low-Rank Experts (MoLE), integrating both shared and separate Low-Rank Adapters (LoRAs). Shared LoRAs capture common feature representations across datasets, while separate LoRAs specialize in diverse generative patterns. A trainable gating mechanism dynamically assigns input tokens to appropriate experts, with a load-balancing loss ensuring uniform expert utilization. The model freezes most CLIP parameters, adapting only LoRA modules and a new MLP classification head with a sigmoid activation. The forward operation in each MLP block is expressed as:
@@ 2.2.4 Mixture of Low-Rank Experts: Liu 2024
2.2.5 MERGING A MIXTURE OF HYPER LORAS: HYPERDET by Cao 2024
Cao et al. Cao et al. (2024) propose a generalizable synthetic image detection framework, HyperDet, leveraging the large pretrained multimodal model, CLIP: ViT-L/14, as the backbone, similar to Liu et al. Liu et al. (2024b), but with few notable innovations. Unlike Liu et al., the authors introduce a novel grouping of Spatial Rich Model (SRM) filters into five distinct groups to generate multiple filtered views of input images, capturing varying levels of pixel artifacts and features. Along this, HyperDet employs Hyper LoRAs, a hypernetwork-based approach that generates Low-Rank Adaptation (LoRA) weights for fine-tuning the CLIP model. These LoRA weights are computed using three types of embeddings: task embeddings, layer embeddings, and position embeddings. The outputs of these LoRA experts are merged to form a unified representation for classification, effectively integrating shared and specific knowledge for generalizable feature extraction. During training, HyperDet fine-tunes the last eight fully connected layers of the CLIP: ViT-L/14 model, along with the newly introduced Hyper LoRAs modules. To address imbalanced optimization, the framework employs a composite binary cross-entropy loss function, incorporating both original and filtered views of the images. This design obtains robust performance in detecting synthetic images across diverse generative models and datasets.
2.2.6 Forgery-Aware Adaptive Transformer (FatFormer) by Liu: 2024
Liu et al. Liu et al. (2024a) introduce FatFormer, a generalizable synthetic image detection framework utilizing the pretrained CLIP model inspired by the work of Ojha et al. Ojha et al. (2023). The authors address the limitations of freezing CLIP’s layers, which hinders the generalization of forgery detection. FatFormer integrates two modules: the Forgery-Aware Adapter (FAA) and Language-Guided Alignment (LGA), for effective adaptation of CLIP’s features. The FAA module extracts forgery artifacts from both image and frequency domains. The Image Forgery Extractor applies lightweight convolution layers to capture low-level artifacts, while the Frequency Forgery Extractor employs Discrete Wavelet Transform (DWT) and grouped attention mechanisms to dynamically aggregate multi-band frequency clues. The final adapted feature representation at each ViT stage is defined as:
@@ 2.2.6 Forgery-Aware Adaptive Transformer (FatFormer) by Liu: 2024
2.2.7 Raising the Bar with CLIP: By Cozzolino 2024
Cozzolino et al. Cozzolino et al. (2024) implement the CLIP: ViT-L/14 pretrained VLM, similar to the approaches in Cao et al. (2024); Liu et al. (2024b); Koutlis and Papadopoulos (2025), for detecting synthetic images with a straightforward yet impactful adjustment. The authors propose generating synthetic images by feeding real-image captions to text-to-image models and then extracting feature vectors for both real and synthetic images using CLIP’s image encoder. Specifically, feature vectors are obtained from the second-to-last layer of the ViT module:
@@ 2.2.7 Raising the Bar with CLIP: By Cozzolino 2024
2.2.8 Representation from Encoder-Decoder for Image Detection by Koutlis: 2025
Koutlis and Papadopoulos Koutlis and Papadopoulos (2025) propose Representations from Intermediate Encoder-Blocks (RINE) to improve synthetic image detection by extracting low-level features from multiple layers of CLIP’s Vision Transformer (ViT). The method captures both low- and high-level visual semantics by concatenating CLS tokens from each intermediate transformer block into a comprehensive feature representation. The extracted CLS tokens from each block $Z_{l}$ are aggregated:
@@ 2.2.8 Representation from Encoder-Decoder for Image Detection by Koutlis: 2025
2.2.9 Towards Explainable Fake Image Detection with Multi-Modal Large Language Models by Ji et al., 2025
Ji et al. Ji et al. (2025a) present a study on explainable AI-generated image detection leveraging Multi-Modal Large Language Models (MLLMs) and prompt engineering. The core idea is to enhance interpretability and robustness by designing six specialized prompt paradigms (P1–P6), each interrogating a distinct visual or logical aspect of an image. These include: (P1) Defect Query for identifying artifacts such as abnormal lighting or unrealistic textures, (P2) Regional Analysis focusing on region-of-interest cues extracted using DINOv2, (P3) Common Sense Reasoning to detect physical or spatial inconsistencies, (P4) Few-Shot Prompting with exemplar comparisons, (P5) Structural Analysis assessing missing or misplaced components, and (P6) Stereotype Matching examining exaggerated or overly uniform features. The outputs of these six paradigms are then fused—either sequentially or via majority voting—to yield a final decision supported by coherent reasoning. The authors constructed a diverse dataset of 2,000 images (1,000 real and 1,000 AI-generated) spanning diffusion, GAN-based, and other generative architectures, ensuring cross-model generalization. They benchmarked four major MLLMs—GPT-4o, GPT-4o-mini, Llama-3.2-Vision-Instruct, and LLaVA-CoT—against advanced but non-reasoning detectors such as AEROBLADE Ricker et al. (2024), CNNSpot Wang et al. (2020b), and so on. The fusion strategy using GPT-4o achieved the highest accuracy (93.4%), outperforming CNNSpot (91.8%) and even the best human annotator (86.3%). The study also revealed that replacing the term “fake” with “generated” in prompts improved model acceptance and reduced rejections, enhancing stability and interpretability.
2.2.10 Interpretable and Reliable Detection via Grounded Reasoning in MLLMs by Ji et al: 2025
Ji et al. Ji et al. (2025b) propose an interpretable and reliable detection framework that advances AI-generated image forensics through grounded multimodal reasoning. The study introduces the FakeXplained dataset, which contains 8,772 high-quality AI-generated images annotated with bounding boxes and descriptive captions highlighting synthesis artifacts such as logical inconsistencies, unnatural textures, and structural errors. Using this dataset, the authors fine-tune the Qwen2.5-VL-32B Bai et al. (2025) model through a two-stage process involving supervised fine-tuning (SFT) followed by Reinforcement Learning from Human Feedback (RLHF) with progressive Group Relative Policy Optimization (GRPO). This progressive training employs structured rewards for classification accuracy, grounding precision, and output format validity, guiding the model to produce coherent and human-aligned reasoning. The resulting model classifies images as real or synthetic as well as localizes flawed regions and provides textual justifications aligned with visual evidence. Furthermore, it demonstrates strong robustness under compression, cropping, and downsampling, as well as generalization to unseen datasets such as FaceForensics++ Rossler et al. (2019) and GPT-4o-generated images. The approach establishes a significant step toward explainable and human-trustworthy AI-generated image detection through grounded reasoning in large multimodal models.
2.2.11 ThinkFake: Reasoning in Multimodal Large Language Models by Huang et al: 2025
The most important aspect of ThinkFake introduced by Huang et al. Huang et al. (2025) is the integration of reasoning into AI-generated image detection. The ThinkFake framework introduces a reasoning-based approach to identifying synthetic imagery using Multimodal Large Language Models (MLLMs/VLMs). Built upon the Qwen2.5-VL-7B model Bai et al. (2025), ThinkFake performs deliberate and interpretable detection through a Forgery Reasoning Prompt that follows a structured chain-of-thought pipeline—quick scan, detailed observation, technical tracing, auxiliary inspection, and final decision—to produce human-readable explanations. The framework incorporates Group Relative Policy Optimization (GRPO) reinforcement learning, where four reward functions—reasoning format, JSON structure, accuracy, and agentic reward—guide structured reasoning and enhance generalization. The agentic reward leverages auxiliary expert agents (UnivFD Ojha et al. (2023), AIDE Yan et al. (2024), and so on) to provide semantic, frequency, and dual-stream feedback for more robust detection. Training follows the GenImage protocol, using 6,000 balanced samples of real and AI-generated images (from ImageNet and Stable Diffusion v1.4, respectively). A small supervised fine-tuning (SFT) set of 638 samples is created using Gemini-1.4-pro for reasoning annotations, while a larger reinforcement learning (RL) set of 5,000 samples is used for GRPO-based training. ThinkFake achieves 84.0% mean accuracy on GenImage and 75.4% on LOKI, surpassing all baseline detectors while maintaining strong interpretability and generalization capability.
2.2.12 ForenX: Towards Explainable AI-Generated Image Detection with MLLMs by Tan et al., 2025
ForenX Tan et al. (2025) proposes an explainable AIGI detector that augments an MLLM (LLaVA-8B Liu et al. (2023) with Llama-3) using a forensic prompt constructed from CLIP-ViT Dosovitskiy (2020); Radford et al. (2021) features and a lightweight forensics projector; the projector creates a forensics embedding supervised by an auxiliary detection loss and mapped into the LLM’s token space, so the LLM consumes three inputs—text tokens, visual content tokens, and the forensic prompt—to produce a yes/no decision with human-readable reasons. The dataset pipeline has two parts: (i) large-scale machine-generated conversations from LLaVA over GenImage and ForenSynths (content Q&A + detection labels), and (ii) ForgReason, a human-aligned set composed of 2,215 realistic Midjourney fakes with box-level artifact descriptions summarized via GPT-4V Achiam et al. (2023), plus 5,000 real and 1,000 fake samples from GenImage to balance fine-tuning. Training proceeds in two stages with LoRA: Stage-1 jointly tunes CLIP-ViT and the MLLM on GenImage/ForenSynths; Stage-2 freezes CLIP-ViT and further instruction-tunes the MLLM on ForgReason to strengthen explanation quality without degrading recognition. ForenX achieves strong cross-source detection (e.g., 97.8% mAcc on GenImage; 94.4% on ForenSynths) and produces grounded, user-study–preferred explanations; it also generalizes to SD-v3 and FLUX images (97.7–97.8% mAcc).
2.2.13 AIGI-Holmes: Towards Explainable and Generalizable AI-Generated Image Detection via MLLMs by Zhou et al., 2025
AIGI-Holmes Zhou et al. (2025) introduces an explainable and generalizable AI-generated image (AIGI) detector that builds on LLaVA-1.6-Mistral-7B with a CLIP-ViT/L-14 visual encoder augmented by an NPR Tan et al. (2024) low-level artifact expert. It constructs the Holmes-Set, a large-scale dataset comprising Holmes-SFTSet (65K images with textual explanations across semantic and artifact-level features) and Holmes-DPOSet (4K contrastive preference pairs refined through expert feedback). The dataset is annotated via a Multi-Expert Jury system combining four MLLMs (Qwen2VL-72B, InternVL2-76B, InternVL2.5-78B, and Pixtral-124B) with structured prompting and human preference refinement. The proposed Holmes Pipeline consists of three training stages: (1) Visual Expert Pre-training using CLIP-ViT and a low-level artifact detector NPR Tan et al. (2024) to establish domain-specific perception; (2) Supervised Fine-Tuning (SFT) to align multimodal representations with textual explanations of real/fake reasoning; and (3) Direct Preference Optimization (DPO) to refine reasoning and align explanations with human judgment. During inference, a Collaborative Decoding strategy fuses predictions from the visual expert and MLLM for robust decision-making. Trained on LLaVA1.6-Mistral-7B AIGI-Holmes achieves 99.2% mean accuracy and 99.9 A.P. on unseen diffusion and autoregressive models (e.g., VAR, FLUX), outperforming prior methods such as RINE, AIDE, and NPR. Its explanations also surpass GPT-4o and Pixtral-124B in BLEU (0.622) and ELO (11.42), demonstrating state-of-the-art interpretability and robustness against perturbations (JPEG, blur, downsampling).
2.2.14 Fake-GPT: Detecting Fake Image via Large Language Model by Fan et al., 2024
Fake-GPT Fan et al. (2024) introduces a paradigm shift in fake image detection by employing a pure Large Language Model (LLM) for AI-generated image detection, without any vision encoders, position embeddings, or multimodal alignment. Instead of visual feature extraction, the authors reformulate the problem as a text-sequence prediction task, converting each RGB image into a serialized textual representation of pixel values (e.g., 32×32 RGB flattened as a string) and fine-tuning an LLM to classify it as real or fake. The proposed system uses Qwen-7B-Chat Bai et al. (2025) as the base model and applies Low-Rank Adaptation (LoRA) fine-tuning to adapt the pre-trained LLM for this unconventional visual task. The model receives a prompt such as “You are a trained fake image detector. Given a string of RGB pixel values, determine if this image is fake.”—allowing the LLM’s sequence reasoning ability to distinguish subtle pixel-level patterns indicative of generative artifacts. For training, real samples were drawn from LSUN, ImageNet, CelebA, CelebA-HQ, and COCO datasets, while fake images were generated using ProGAN and evaluated across StyleGAN, StyleGAN2, BigGAN, CycleGAN, StarGAN, GauGAN, and DeepFake generators. All inputs were resized to 32×32 to fit within the token limit, showcasing the model’s robustness under low-resolution settings. Experiments demonstrated competitive accuracy—88.2% mean across cross-model evaluations and up to 97.4% on CycleGAN images—outperforming several CNN-based baselines like BiHPF, LGrad, and FrePGAN. When extended to diffusion-generated datasets, Fake-GPT maintained a mean accuracy of 83.8%, confirming strong cross-model generalization. In summary, Fake-GPT pioneers the application of pure LLMs for image authenticity detection, reframing fake image detection as a language modeling task. It eliminates vision-specific modules, simplifies training, and demonstrates that large-scale sequence models can serve as universal detectors for AI-generated imagery across both GAN and diffusion domains.
Comparative Analysis of Vision-Language Models
The earlier generation of vision-language approaches focuses primarily on leveraging pretrained VLMs such as CLIP to improve the generalization of AI-generated image detection through robust visual–textual feature alignment. Ojha et al. Ojha et al. (2023) initiated this direction by mapping images into CLIP’s semantic space to enable cross-model detection, while Wu et al. Wu et al. (2023) extended it with language-guided textual prompts to refine semantic supervision. Subsequent works, including Zhu et al. Zhu et al. (2023), Liu et al. Liu et al. (2024b), and Cao et al. Cao et al. (2024), progressively enhanced adaptability through feature-level regularization, low-rank adapters, and hypernetwork-driven parameterization, thereby strengthening cross-generator robustness. FatFormer Liu et al. (2024a) advanced this paradigm by incorporating frequency-domain features and language-guided alignment, while Cozzolino et al. Cozzolino et al. (2024) simplified CLIP-based detection through a linear SVM classifier to isolate higher-level semantics less sensitive to low-level perturbations. Finally, Koutlis and Papadopoulos Koutlis and Papadopoulos (2025) further optimized representation learning by aggregating intermediate CLIP features to capture both global and local semantics efficiently. Collectively, these studies form a coherent progression—from feature-space distance learning to parameter-efficient fine-tuning and adaptive multimodal integration—marking a shift from handcrafted forensic cues toward transferable, foundation-model-based visual-text alignment for synthetic image detection.
Comparative Analysis of MLLM Models
All studies converge toward a shared goal of enhancing the interpretability, reliability, and generalizability of AI-generated image detection through language-based reasoning frameworks. The earliest efforts (from 2024 to 2025), such as Fake-GPT Fan et al. (2024) and ForenX Tan et al. (2025), establish two divergent yet complementary directions—Fake-GPT reformulates image forensics as a pure sequence modeling task using only a textual representation of pixels, while ForenX introduces the first multimodal prompting mechanism that explicitly injects forensic embeddings into an MLLM for explainable decisions. Ji et al. Ji et al. (2025a, b) advance the field with Towards Explainable Detection and Grounded Reasoning, which evolve from handcrafted forensic prompts to structured, grounded reasoning guided by reinforcement learning and visual grounding, deepening interpretability and human alignment. ThinkFake Huang et al. (2025) continues this trajectory by explicitly modeling a reasoning chain-of-thought through GRPO optimization and multi-agent feedback, bridging MLLM interpretability with robustness. Finally, AIGI-Holmes Bai et al. (2025) represents the culmination of these trends, integrating multi-expert jury annotation, large-scale instruction tuning, and collaborative decoding to achieve both explainability and cross-domain generalization. Collectively, these works demonstrate a clear methodological progression—from textual reformulation (Fake-GPT), to multimodal prompting (ForenX), to reasoning-driven and human-aligned explanation systems (Ji et al., ThinkFake), culminating in a comprehensive, expert-supervised reasoning pipeline (AIGI-Holmes). The principal differences thus lie in their modality design (pure LLM vs. MLLM), training paradigm (prompt engineering vs. SFT/RLHF/DPO), and explanation grounding (textual, visual, or collaborative), marking a coherent evolution toward explainable and reasoning-based forensic detection.